OpenStudy (anonymous):

What is a line equation of a line perpendicular to line y=-(3/4)x + 4?

OpenStudy (anonymous):

find the slope of y first

OpenStudy (anonymous):

The Perpendicular line has a slope opposite to the slope of the origional line, basically the inverse.

OpenStudy (anonymous):

the slope of y * the slope of another line = -1

OpenStudy (anonymous):

The Perpendicular line has a slope opposite to the slope of the origional line, basically the inverse.

OpenStudy (anonymous):



OpenStudy (anonymous):

So basically y= 4/3x ?

OpenStudy (anonymous):

Ohh ok. Thank you all.

OpenStudy (anonymous):

No, that's wrong!

OpenStudy (anonymous):

What would it be?

OpenStudy (anonymous):

Y= 3/4 ?

OpenStudy (anonymous):

Y = 4/3 if you're in a Perpendicular line, parallel lines have the same slope.